Dear math teacher,
I am having the following difficulties with this problem:
"A box contains 7 red cards, 6 white cards and 4 blue cards. How many selections of three cards can be made so that a) all three are red, b) none are red?"
I solved the problem correctly but then I tried to solve it the second way - the shorter and less time consuming way. I was supposed to get the same result but did not. Please let me know, why I did not get the same answer.
Here is what I did by Method 1 for part a):
R W B
n 7 6 4
r 3
7C3 = 35 selections
Here is what I did by Method 1 for part b):
R W B
n 7 6 4
r n = 6+4 = 10
r = 3
10C3 = 10!/(10-3)! = 120 selections.
Method 2 part b)
n (total R, W, B) = 7+6+4 = 17
r = 3
17C3 - 7C3 = 680-35 = 645 selections and this answer does not match the answer obtained by Method 1 for part b). Would you tell me why? Where did I do a mistep? Here is how reasoned it out: Total Number of Ways - Number of Ways to Pull Red Cards = Number of Ways to Get Other Colors
Why would that be wrong to think like that?
Thank you very much for your help.
Respectfully,

Ivanka

Answer by stanbon(75887)

"A box contains 7 red cards, 6 white cards and 4 blue cards. How many selections of three cards can be made so that
a) all three are red,
Easiest Way:
# of ways to get 3 red: 7C3 = (7*6*5)/(1*2*3) = 35
------------------------
b) none are red?
Easiest Way:
# of ways to get 3 not red: 10C3 = (10*9*8)/(1*2*3) = 120
--
Your Question::::::
17C3 - 7C3 = 680-35 = 645
The 645 includes all the combinations that allow
1 red in the group of 3 or 2 reds in the group of 3.
---
Your 35 is the # of groups with ALL red
Your 120 is the # of groups with NO red
============================================
Cheers,
Stan H.